A bystander notes that the coordinates of an incident are (10 m, 0 m, 0 m, 7.0 s). What are the

coordinates of this accident as seen by the reporter traveling at 6.0 m/s towards the accident?

Assume that at t = 0, the reporter and by bystander were at the same position.

near Wolofonward at 1 mis polotive to the hue Low


1
Expert's answer
2021-06-25T12:05:01-0400

Minutes to seconds

"(10 m, 0 m, 0 m, 7.0 s) = (10 \times 60+ 0\times 60+0 \times 60+7.0 s) =600+7=607 s"

Therefore, "Distance = Time \times Speed = 607 s \times 6.0 m/s = 3642 m"

The reporter was 3642 m away from the accident.